can someone help me solve this problem?
1. consider the function f(x)=x^2 + 6x - 2
A. find h, the x-coordinate of the vertex of this parabola.
answer_______________________________________________
B. Substitute the two whole number values immediately to the left and right of h into the function to find the cooresponding y. Fill in the table below. Make sure your x-values are in increasing order in your table.
x y
__h=___ ________
_____ ________
_____ ________
_____ ________
_____ ________
C. Use microsoft exel to graph the function by plotting the points found in the table in part B.
need help ASAP
THANKS

Answer by scott8148(6628)

A) the vertex lies on the axis of symmetry ___ the equation is x=-b/2a ___ x=-6/(2*1) ___ x=-3 ___ this is the h value

B) the two numbers to the left (less than) of h are -5 and -4
___ the two to the right (greater than) are -2 and -1
___ plug the numbers as x-values into f(x) and find the corresponding y-values ___ y IS f(x); same thing, different name
___ enter the values in the table

C) good luck with excel